OpenStudy (anonymous):

Find the volume generated by rotating the region enclosed by the graphs of f(x)=x^2 & g(x)=x about the y-axis.

OpenStudy (sleepyhead314):

when you see something like "rotated about the y axis" get the equations you were given in terms of y so y = x^2 --> x = ? and y = x --> x = ?

OpenStudy (anonymous):

that's all i have sleepy head

OpenStudy (sleepyhead314):

? you mean you don't know how to get it in terms of y? just solve for x

OpenStudy (anonymous):

really?

OpenStudy (sleepyhead314):

yep :P

OpenStudy (anonymous):

x^2 = sqrt(y), -sqrt(y) x = y

OpenStudy (sleepyhead314):

yep :) and find the point where they intersect

OpenStudy (anonymous):

lol :o let me try

OpenStudy (anonymous):

(0,0) :S

OpenStudy (sleepyhead314):

and... ? :P

OpenStudy (anonymous):

hmm... make those the min and max right?

OpenStudy (sleepyhead314):

lol there's another point I mean :P

OpenStudy (anonymous):

lol oops. shoud I add the originals?

OpenStudy (anonymous):

oh (1,1)

OpenStudy (sleepyhead314):

yeah lol ;) and if you rotate it, you'll get a washer >.> so pi times the integral from 0 to 1 of y minus y^2 dy the 0 came from (0,0) the 1 came from (1,1) the y was from (root y)^2 = y the y^2 was from (y)^2 = y^2 :P

OpenStudy (anonymous):

so min = 0 and max = 1

OpenStudy (sleepyhead314):

yeah :)

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

-2 pie / 15?

OpenStudy (sleepyhead314):

hmmm that's not quite what I got...

OpenStudy (anonymous):

:( darn

OpenStudy (anonymous):

2pi/15?

jigglypuff314 (jigglypuff314):

erm... integral of y-y^2 --> (1/2)y^2 - (1/3)y^3

OpenStudy (anonymous):

hey it's jiggly puff!! k one sec.

OpenStudy (anonymous):

i got a big number. I took "y-y^2" as y to y^2

OpenStudy (sleepyhead314):

huh? y-y^2 as in y minus y squared

OpenStudy (anonymous):

lol my brain hurts one sec

OpenStudy (anonymous):

9/64

OpenStudy (sleepyhead314):

>.> if you just plug in 1 into (1/2)y^2 - (1/3)y^3 http://www.wolframalpha.com/input/?i=pi+times+the+integral+from+0+to+1+of+y+minus+y%5E2+dy

OpenStudy (anonymous):

pi/6??

OpenStudy (sleepyhead314):

yep :)

OpenStudy (anonymous):

<3